Write an equation of the line that passes through the given point (0,6) and is parallel to 7y-35x=4. Write the answer in slope intercept form:
a) y=5x+6
b) y=5x-6
c) -1/6x+4
d) -1/5x+6
e) y=-1/5x+4
explain your answer

Answer by stanbon(75887)

Write an equation of the line that passes through the given point (0,6) and is parallel to 7y-35x=4. Write the answer in slope intercept form.
----------
Form of the equation is y = mx + b
You need to find m and b.
--------
The given equation can be written in the form y = (35/7)x + (4/7)
So the slope is (35/7) = 5
---
You know y = 6 when x = 0 and you know m = 5, so solve for "b"
6 = 5*0 + b
b = 6
-------
Equation:
y = 5x + 6
====================
Cheers,
Stan H.
====================